A boy is 17 years old and his sister is twice as old. When the boy is 23 what will be the age of his sister.
I would like for you to show me how you got the answer also because I have a whole worksheet trying to solve people age.

Answer by stanbon(75887)

A boy is 17 years old and his sister is twice as old. When the boy is 23 what will be the age of his sister.
----------
Equations:
b = 17; s = 2*17 = 34
----
The boy will be 23 in 6 years.
In 6 years his sister will be 34+6 = 40.
------------------------
Cheers,
Stan H.